OpenStudy (anonymous):

Will someone please help me? Factor completely: -7x^2+14x-21

OpenStudy (anonymous):

-7(x^2-2x+3)

OpenStudy (anonymous):

Okay you. Would you mind explaining how you did that?

OpenStudy (anonymous):

haha ok well you can check the work by simply factoring it out again for example (-7)(x^2)=-7^2 and (-7) (-2x)=14x then (-7)(3)=-21

OpenStudy (anonymous):

Okay that makes sense :) thank you!!

OpenStudy (anonymous):

no prob.